# Stock Movement Prediction with Multimodal Stable Fusion via Gated Cross-Attention Mechanism

### v1

@@ Abstract
keywords:
Stock movement prediction, Multimodal fusion, Gated cross-attention, Financial forecasting
1 Introduction
With the continual increase in the capitalization of the stock market, trading stocks has become an attractive investment instrument for many investors. The task of predicting stock movements, which focuses on forecasting future trends of a stock’s prices, benefits the right selection of stocks and is of great significance for investment decisions [1]. Stock movement prediction is a very difficult task. Traditional stock movement prediction methods apply machine learning techniques to perform data mining on financial indicators [2, 3], social media documents [4, 5], or market relationships [1, 9]. These methods employ a specific neural network encoder, which is suitable for a particular type of information, to represent stock features.
However, the price trend of a stock could be influenced by a variety of factors, and the extent of these factors’ impacts can differ. These factors, collected from various data sources, are often presented in different modalities, including numerical indicators (e.g., prices, financial indicators), textual content (e.g., public opinions, news) and graph structures (e.g., industry relations, investment relations). How to effectively utilize the features of these multimodal factors to accurately predict stock price movements has always been a challenge. Recent multimodal approaches try to solve the stock movement prediction task considering two sources of information, such as the integration of the price and the text signal [10, 11], or a blend of social media text and a correlation graph [12, 13]. These methods mostly apply multiple specific encoders subsequently for each modality and integrate features with simplified fusion strategies such as vector concatenation and conventional attention mechanism, without considering distinctions and conflicts between modalities.
Meanwhile, there is still a distance between the task objectives and the actual demands. The latest multimodal deep learning methods focused only on predicting the rise and fall of stock prices, which is considered a binary classification problem [14, 15, 16]. However, slight fluctuation of stock prices is also an important phenomenon in real scenarios, usually called a sideways trend or consolidation, indicating that the forces of supply and demand are nearly equal before a new trend  [^1] https://www.investopedia.com/terms/s/sidewaystrend.asp. However, some early studies involve multi-label classification with traditional deep learning models such as Long Short-Term Memory (LSTM) [17] and statistical methods such as random forests [18], there is still a research gap between multimodal fusion and multi-label stock prediction. Unifying these two settings has an unquestionably practical meaning for stock-trading scenarios.
@@ 1 Introduction
Challenge 1 involves the use of overly simplistic methods to tackle complex prediction tasks in current research. This includes the introduction of too few modalities and the application of basic fusion techniques, which fail to account for the diversity inherent in multimodal prediction. To predict stock movements, analysts typically use continuous financial indicators (such as opening and closing prices), textual information from various sources (including news, tweets, and blogs), and graph data reflecting market relationships (such as industry sectors and supply chains). However, existing research on stock movement prediction has been limited to unimodal [2, 3, 5] or bimodal [15, 19, 20] analysis, resulting in predictions that often diverge significantly from actual market trends. Furthermore, multimodal fusion techniques significantly impact overall predictive accuracy. Current methods typically employ single-modal encoding followed by multimodal integration [21], often defaulting to simple vector concatenation for fusion [22, 23, 24], which fails to facilitate interactivity between different data types. Other studies employ co-attention mechanisms to fuse multimodal data [25, 26], yet often overlook the distinct contributions of each modality to the specific task. All these shortcomings of existing research drive us to conduct further studies on predicting stock trends with more data sources using a more optimized and efficient model.
@@
Challenge 2 involves a data inconsistency problem while utilizing multiple data sources to address the prediction task. This includes data sparsity and semantic conflicts, which markedly impact the precision of predictions. As shown in Fig. 1, the data we gather from various sources frequently contains inconsistencies due to missing values, hindering temporal alignment across all modalities. Furthermore, semantic conflicts may arise when different modalities convey contradictory meanings. Traditional fusion techniques applied to these data sources can compromise or diminish predictive accuracy. Previous studies on multimodal stock movement prediction have simplified the issue by incorporating fewer sources of information [14, 16, 27, 28] or have neglected it altogether [29, 30]. To address the problem of data sparsity among modalities, most previous methods concentrate on filtering and smoothing in the presence of missing values to handle uncertainty by a factorized inference model [31] or a heterogeneous graph model [32]. Another method [33] investigates that Transformer models are sensitive to incomplete modalities and to automatically search for an optimal strategy. Meanwhile, very few studies have focused on the problem of semantic conflicts between modalities. One way is to identify noise with a density estimation block based on the inherent correlation between modalities [34]. Another work addresses the problem of the imbalance contribution of data sources by learning a network to determine the level of noise of each encoder [35]. Although these studies improve the performance of multimodal tasks by handling particular data problems between modalities, they cannot be universally applied to address this challenge and also increase the complexity of the overall architecture. This urges us to come up with a more general approach to solve these unstable fusion problems for stock movement prediction.
@@
  • We implement, for the first time, a framework to solve multimodal fine-grained stock movement prediction, named MSGCA, which efficiently utilizes features from various modalities. This framework characterizes and integrates the indicator sequences of stock prices, the dynamic documents of sentiment and news, and the relational graph of industry sectors, to predict the up, flat, and down trends for stocks.
@@
The remainder of the paper is organized as follows. First, we review the related work and make a brief comparison between our method and previous studies in Section 2. Then we provide the preliminary and task formulation of our study in Section 3. The implementation of our MSGCA stock movement prediction method is presented in Section 4. Section 5 shows the experiments and analysis, followed by the conclusion and future work in Section 6.
2 Literature review
2.1 Stock movement prediction
@@
Stock movement prediction is challenging due to the market being highly stochastic and temporally dependent on predictions from chaotic data [10]. We summarize the existing stock movement prediction methods in Table 1. Traditional methods focus on utilizing single-modal input with statistical [4, 18, 7] or simple deep learning methods [1, 2, 3]. For example, an early work named JST [4] extracts topics and sentiments simultaneously using a linear kernel SVM and utilizes them for stock market prediction. MFNN [3] learns an end-to-end model with integration of convolutional and recurrent neurons for feature extraction on financial time series samples. IDTLA [7] proposes an instance-based deep transfer learning models with attention mechanism to address the lack of historical stock data. CT-GCNN [6] constructs a conceptual-temporal graph and applies a graph convolutional neural network model to map the linkage effect and predict the stock price movement. However, all of the above methods focus on a singular modality in one of indicators, text, or graph, thereby providing limited information for the multifaceted task. In addition, these works utilize LSTM or CNN variants that are limited in their representational capabilities. To address these problems, recent efforts on stock movement prediction focus on introducing more types of information [10, 11, 30] and using modern deep learning methods such as the attention mechanism [9, 16, 27, 36, 37, 8] and large language models [14, 19]. For example, TRAN [9] proposes a time-aware relational attention network for graph-based stock movement prediction to capture time-varying correlation strengths between stocks. MAC [11] proposes a multisource aggregated method that incorporates the numerical features, market-driven news, and the sentiments of their related stocks. ALSTM [37] complete the task using components that include feature mapping, LSTM, temporal attention, and prediction layer. SLOT [16] learns the unified embeddings of prices and tweets from self-supervised learning and uses ALSTM for prediction.
@@
All the aforementioned methods focus on predicting the up and down trends of stock prices given various stock features. However, the flat trend of a stock is also important for prediction, which indicates that the forces of demand are nearly equal before a new period. Some previous studies aim to solve a fine-grained prediction but with single-modal input. In contrast, our proposed MSGCA tries to perform a stable multimodal fusion by introducing three types of modalities (indicator, documents, and graph) to solve the task with labels of up, flat and down, which fills the gap for fine-grained prediction with multimodal features as shown in Table 1.
Table 1: Summarization of existing stock movement prediction methods
@@
Coarse-grained Prediction JST [4] MFNN [3] CT-GCNN [6] IDTLA [7] ALSTM [37] DTML [36] TRAN [9] StockNet [10] MAC [11] AStock [19] ESTIMATE [27] SLOT [16]
@@
2.2 Multimodal fusion strategies
Multimodal fusion is the process of combining collected data from various modalities for analysis tasks [40], and the fusion of multimodal data can provide surplus information with an increase in accuracy of the overall result [41]. Most early multimodal fusion strategies implement the framework of encoder and concatenation process, which focuses on using a vectorized concatenation operator to generate the final features [22, 23, 24]. These simple fusion methods overlook the interference of noisy information while performing useful feature extraction [22]. Some other works utilize the attention and co-attention mechanism to simultaneously integrate information from all modalities [25, 42, 43], but ignore the differences between modalities for the targeted tasks. These methods involve on filtering noise while jointly learning the multimodal features. Modern approaches focus on more sufficient multimodal interactions [26, 44]. These studies employ self-attention and co-attention for intro-modal and inter-modal interaction, respectively, but suffer from using redundant information and architecture during the process, and further impair the efficiency of the framework [45].
@@ 2.2 Multimodal fusion strategies
2.3 Cross- and gated-attention mechanisms
A powerful and robust model to represent features from multiple input sequences requires components that can effectively aggregate information relevant to the task from each source, and cross-attention is considered as an effective mechanism for feature integration [46]. Existing cross-attention studies treat one sequence as query, and the other sequences as key and value to perform the encoder [47, 48, 49, 50]. For example, IRENE [49] fuses textual and visual features with two bidirectional cross-attention blocks to achieve mutual integration between two types of input. Another work [46] implements a tri-modal cross-attention architecture to perform cross-attention with combinatorial input from three sources. In addition, the gating mechanism has been shown to be essential for recurrent neural networks (RNNs) to achieve state-of-the-art performance [51]. Gated linear unit (GLU) can reduce the problem of gradient vanishing by providing a linear path for gradients while maintaining non-linear capabilities [52]. Gated attention unit to formulate attention and GLU as a unified layer, which allows the use of a much simpler attention mechanism and results in higher computational efficiency [53].
@@ 3 Definition and formulation
  • Relational graph, shows the intra-industry relationships between companies [38, 54], denoted as $\mathcal{G}=\{\mathcal{E},\mathcal{R},\mathcal{U}\}$, where $\mathcal{E}$ is the set of entities and $\mathcal{R}$ is the set of relations. $\mathcal{U}=\{(e_{s},r,e_{d})|e_{s},e_{d}\in\mathcal{E},r\in\mathcal{R}\}$ is the set of triplets in the graph $\mathcal{G}$, where $e_{s}$, $e_{d}$ and $r$ are the start entity, the end entity, and the relation, respectively.
@@
4 The method of MSGCA
4.1 Overall framework
The framework of our proposed method MSGCA is shown in Fig. 2. The procedure is structured into three phases, adhering to an encoder-fusion-decoder architecture. During the multimodal encoding phase, projections from three modalities are transposed into a cohesive latent space through employing various representational methods as encoders. The gated cross-feature fusion phase implements a stable multimodal integration strategy to synthesize a unified feature representation through a pair of gated cross-attention networks. The movement prediction phase compresses both temporal and feature dimensions, aligning them with the granularity of the trend labels, thereby enabling precise prediction of stock movement at timestamp $T+1$. In Table 2, we summarize the notations used in our framework.
@@ 4.1 Overall framework
Table 2: Notations used in our method
@@
4.2 Multimodal source encoding
Analysis of stock movement is based on the idea that historical market behavior can inform traders about potential future trends. Conventional methods for stock movement prediction focus on employing statistical models on a range of financial indicators, including open prices, highest prices, lowest prices, closed prices, and trading volumes [2, 4, 18]. Recent deep learning approaches are apt to introduce more types of data sources, such as tweet text, and market relational data [9, 16, 38]. The above methods encompass the use of indicators, documents, and graphs, which inspires us to concentrate our efforts on the management of these three distinct data sources. We implement a trimodal encoding module as the first phase of MSGCA to concurrently encode each type of multimodal information into a unified latent dimensional space. Each encoder is described as follows.
@@ 4.2 Multimodal source encoding
$\textbf{v}_{c}=\textbf{s}_{c}\textbf{W}_{ic}+b_{ic},\textbf{v}_{o}=\textbf{s}_{o}\textbf{W}_{io}+b_{io},\textbf{v}_{h}=\textbf{s}_{h}\textbf{W}_{ih}+b_{ih}$ (2)
@@
4.3 Gated cross-feature fusion
@@
$\textbf{H}_{a}=\textbf{H}_{i,d}^{l}\textbf{W}_{a}+b,\textbf{H}_{b}=Sigmoid(\textbf{H}_{i}\textbf{W}_{b}+b^{{}^{\prime}})$ (11)
@@
$\textbf{H}_{a}=\textbf{H}_{i,d,g}^{l}\textbf{W}_{a}+b,\textbf{H}_{b}=Sigmoid(\textbf{H}_{i,d}\textbf{W}_{b}+c)$ (15)
@@
4.4 Fine-grained movement prediction

### v2

@@ Abstract
Index Terms:
Stock movement prediction, Multimodal fusion, Gated cross-attention
I Introduction
With the continual increase in the capitalization of the stock market, trading stocks has become an attractive investment instrument for many investors. The task of predicting stock movements, which focuses on forecasting future trends of a stock’s prices, benefits the right selection of stocks and is of great significance for investment decisions [1]. Stock movement prediction is a very difficult task. Traditional stock movement prediction methods apply machine learning techniques to perform data mining on financial indicators [2] [3], social media documents [4][5], or market relationships [1] [6]. These methods employ a specific neural network encoder, which is suitable for a particular type of information, to represent stock features.
However, the price trend of a stock could be influenced by a variety of factors, and the extent of these factors’ impacts can differ. These factors, collected from various data sources, are often presented in different modalities, including numerical indicators (e.g., prices, financial indicators), textual content (e.g., public opinions, news) and graph structures (e.g., industry relations, investment relations). How to effectively utilize the features of these multimodal factors to accurately predict stock price movements has always been a challenge. Recent multimodal approaches try to solve the stock movement prediction task considering two sources of information, such as the integration of the price and the text signal [7] [8], or a blend of social media text and a correlation graph [9] [10]. These methods mostly apply multiple specific encoders subsequently for each modality and integrate features with simplified fusion strategies such as vector concatenation and conventional attention mechanism, without considering distinctions and conflicts between modalities.
Meanwhile, there is still a distance between the task objectives and the actual demands. The latest multimodal deep learning methods focused only on predicting the rise and fall of stock prices, which is considered a binary classification problem [11] [12] [13]. However, slight fluctuation of stock prices is also an important phenomenon in real scenarios, usually called a sideways trend or consolidation, indicating that the forces of supply and demand are nearly equal before a new trend  [^1] https://www.investopedia.com/terms/s/sidewaystrend.asp. However, some early studies involve multi-label classification with traditional deep learning models such as Long Short-Term Memory (LSTM) [14] and statistical methods such as random forests [15], there is still a research gap between multimodal fusion and multi-label stock prediction. Unifying these two settings has an unquestionably practical meaning for stock-trading scenarios.
@@ I Introduction
Challenge 1 involves the use of overly simplistic methods to tackle complex prediction tasks in current research. This includes the introduction of too few modalities and the application of basic fusion techniques, which fail to account for the diversity inherent in multimodal prediction. To predict stock movements, analysts typically use continuous financial indicators (such as opening and closing prices), textual information from various sources (including news, tweets, and blogs), and graph data reflecting market relationships (such as industry sectors and supply chains). However, existing research on stock movement prediction has been limited to unimodal [2] [3] [5] or bimodal [12] [16] [17] analysis, resulting in predictions that often diverge significantly from actual market trends. Furthermore, multimodal fusion techniques significantly impact overall predictive accuracy. Current methods typically employ single-modal encoding followed by multimodal integration [18], often defaulting to simple vector concatenation for fusion [19] [20] [21], which fails to facilitate interactivity between different data types. Other studies employ co-attention mechanisms to fuse multimodal data [22] [23], yet often overlook the distinct contributions of each modality to the specific task. All these shortcomings of existing research drive us to conduct further studies on predicting stock trends with more data sources using a more optimized and efficient model.
@@
Challenge 2 involves a data inconsistency problem while utilizing multiple data sources to address the prediction task. This includes data sparsity and semantic conflicts, which markedly impact the precision of predictions. As shown in Fig. 1, the data we gather from various sources frequently contains inconsistencies due to missing values, hindering temporal alignment across all modalities. Furthermore, semantic conflicts may arise when different modalities convey contradictory meanings. Traditional fusion techniques applied to these data sources can compromise or diminish predictive accuracy. Previous studies on multimodal stock movement prediction have simplified the issue by incorporating fewer sources of information [11] [13] [24] [25] or have neglected it altogether [26] [27]. To address the problem of data sparsity among modalities, most previous methods concentrate on filtering and smoothing in the presence of missing values to handle uncertainty by a factorized inference model [28] or a heterogeneous graph model [29]. Another method [30] investigates that Transformer models are sensitive to incomplete modalities and to automatically search for an optimal strategy. Meanwhile, very few studies have focused on the problem of semantic conflicts between modalities. One way is to identify noise with a density estimation block based on the inherent correlation between modalities [31]. Another work addresses the problem of the imbalance contribution of data sources by learning a network to determine the level of noise of each encoder [32]. Although these studies improve the performance of multimodal tasks by handling particular data problems between modalities, they cannot be universally applied to address this challenge and also increase the complexity of the overall architecture. This urges us to come up with a more general approach to solve these unstable fusion problems for stock movement prediction.
@@
  • We implement, for the first time, a framework to solve multimodal fine-grained stock movement prediction, named MSGCA [^2] The code of MSGCA and the data for training and evaluation are available at: https://github.com/changzong/MSGCA., which efficiently utilizes features from various modalities. This framework characterizes and integrates the indicator sequences of stock prices, the dynamic documents of sentiment and news, and the relational graph of industry sectors, to predict the up, flat, and down trends for stocks.
@@
The remainder of the paper is organized as follows. First, we review the related work and make a brief comparison between our method and previous studies in Section II. Then we provide the preliminary and task formulation of our study in Section III. The implementation of our MSGCA stock movement prediction method is presented in IV. Section V shows the experiments and analysis, followed by the conclusion and future work in Section VI.
II Related Work
II-A Stock Movement Prediction
@@
Stock movement prediction is challenging due to the market being highly stochastic and temporally dependent on predictions from chaotic data [7]. We summarize the existing stock movement prediction methods in Table I. Traditional methods focus on utilizing single-modal input with statistical [4] [15] or simple deep learning methods [1] [2] [3]. For example, an early work named JST [4] extracts topics and sentiments simultaneously using a linear kernel SVM and utilizes them for stock market prediction. MFNN [3] learns an end-to-end model with integration of convolutional and recurrent neurons for feature extraction on financial time series samples. TGC [1] proposes a temporal graph convolution, which jointly models the temporal evolution and the relation network of stocks. However, all of the above methods focus on a singular modality in one of indicators, text, or graph, thereby providing limited information for the multifaceted task. In addition, these works utilize LSTM or CNN variants that are limited in their representational capabilities. To address these problems, recent efforts on stock movement prediction focus on introducing more types of information [7] [8] [27] and using modern deep learning methods such as the attention mechanism [6] [13] [24] [33] [34] [35] and large language models [11] [16]. For example, TRAN [6] proposes a time-aware relational attention network for graph-based stock movement prediction to capture time-varying correlation strengths between stocks. MAC [8] proposes a multisource aggregated method that incorporates the numerical features, market-driven news, and the sentiments of their related stocks. ALSTM [34] complete the task using components that include feature mapping, LSTM, temporal attention, and prediction layer. SLOT [13] learns the unified embeddings of prices and tweets from self-supervised learning and uses ALSTM for prediction.
@@
All the aforementioned methods focus on predicting the up and down trends of stock prices given various stock features. However, the flat trend of a stock is also important for prediction, which indicates that the forces of demand are nearly equal before a new period. Some previous studies aim to solve a fine-grained prediction but with single-modal input. In contrast, our proposed MSGCA tries to perform a stable multimodal fusion by introducing three types of modalities (indicator, documents, and graph) to solve the task with labels of up, flat and down, which fills the gap for fine-grained prediction with multimodal features as shown in Table I.
TABLE I: Summarization of Existing Stock Movement Prediction Methods.
@@
Coarse-grained Prediction JST [4] CapTE [5] MFNN [3] ALSTM [34] DTML [33] TRAN [6] StockNet [7] MAC [8] AStock [16] ESTIMATE [24] SLOT [13]
@@
II-B Multimodal Fusion Strategies
Multimodal fusion is the process of combining collected data from various modalities for analysis tasks [37], and the fusion of multimodal data can provide surplus information with an increase in accuracy of the overall result [38]. Most early multimodal fusion strategies implement the framework of encoder and concatenation process, which focuses on using a vectorized concatenation operator to generate the final features [19] [20] [21]. These simple fusion methods overlook the interference of noisy information while performing useful feature extraction [39]. Some other works utilize the attention and co-attention mechanism to simultaneously integrate information from all modalities [22] [40] [41], but ignore the differences between modalities for the targeted tasks. These methods involve on filtering noise while jointly learning the multimodal features. Modern approaches focus on more sufficient multimodal interactions [23] [42]. These studies employ self-attention and co-attention for intro-modal and inter-modal interaction, respectively, but suffer from using redundant information and architecture during the process, and further impair the efficiency of the framework [43].
@@ II-B Multimodal Fusion Strategies
II-C Cross- and Gated-Attention Mechanisms
A powerful and robust model to represent features from multiple input sequences requires components that can effectively aggregate information relevant to the task from each source, and cross-attention is considered as an effective mechanism for feature integration [44]. Existing cross-attention studies treat one sequence as query, and the other sequences as key and value to perform the encoder [45] [46] [47] [48]. For example, IRENE [47] fuses textual and visual features with two bidirectional cross-attention blocks to achieve mutual integration between two types of input. Another work [44] implements a tri-modal cross-attention architecture to perform cross-attention with combinatorial input from three sources. In addition, the gating mechanism has been shown to be essential for recurrent neural networks (RNNs) to achieve state-of-the-art performance [49]. Gated linear unit (GLU) can reduce the problem of gradient vanishing by providing a linear path for gradients while maintaining non-linear capabilities [50]. Gated attention unit to formulate attention and GLU as a unified layer, which allows the use of a much simpler attention mechanism and results in higher computational efficiency [51].
@@ III Definition and formulation
  • Relational graph, shows the intra-industry relationships between companies [35] [52], denoted as $\mathcal{G}=\{\mathcal{E},\mathcal{R},\mathcal{U}\}$, where $\mathcal{E}$ is the set of entities and $\mathcal{R}$ is the set of relations. $\mathcal{U}=\{(e_{s},r,e_{d})|e_{s},e_{d}\in\mathcal{E},r\in\mathcal{R}\}$ is the set of triplets in the graph $\mathcal{G}$, where $e_{s}$, $e_{d}$ and $r$ are the start entity, the end entity, and the relation, respectively.
@@
IV Methodology
IV-A Framework of MSGCA
The framework of our proposed method MSGCA is shown in Fig. 2. The procedure is structured into three phases, adhering to an encoder-fusion-decoder architecture. During the multimodal encoding phase, projections from three modalities are transposed into a cohesive latent space through employing various representational methods as encoders. The gated cross-feature fusion phase implements a stable multimodal integration strategy to synthesize a unified feature representation through a pair of gated cross-attention networks. The movement prediction phase compresses both temporal and feature dimensions, aligning them with the granularity of the trend labels, thereby enabling precise prediction of stock movement at timestamp $T+1$. In Table II, we summarize the notations used in our framework.
@@ IV-A Framework of MSGCA
TABLE II: Notations Used in Our Method.
@@
IV-B Multimodal Source Encoding
Analysis of stock movement is based on the idea that historical market behavior can inform traders about potential future trends. Conventional methods for stock movement prediction focus on employing statistical models on a range of financial indicators, including open prices, highest prices, lowest prices, closed prices, and trading volumes [2] [4] [15]. Recent deep learning approaches are apt to introduce more types of data sources, such as tweet text, and market relational data [6] [13] [35]. The above methods encompass the use of indicators, documents, and graphs, which inspires us to concentrate our efforts on the management of these three distinct data sources. We implement a trimodal encoding module as the first phase of MSGCA to concurrently encode each type of multimodal information into a unified latent dimensional space. Each encoder is described as follows.
@@ IV-B Multimodal Source Encoding
$\textbf{v}_{c}=\textbf{s}_{c}\textbf{W}_{ic}+b_{ic}$ (2)
$\textbf{v}_{o}=\textbf{s}_{o}\textbf{W}_{io}+b_{io}$ (3)
$\textbf{v}_{h}=\textbf{s}_{h}\textbf{W}_{ih}+b_{ih}$ (4)
@@
IV-C Gated Cross-Feature Fusion
@@
$\textbf{H}_{a}=\textbf{H}_{i,d}^{l}\textbf{W}_{a}+b$ (13)
$\textbf{H}_{b}=Sigmoid(\textbf{H}_{i}\textbf{W}_{b}+b^{{}^{\prime}})$ (14)
@@
$\textbf{H}_{a}=\textbf{H}_{i,d,g}^{l}\textbf{W}_{a}+b,$ (18)
$\textbf{H}_{b}=Sigmoid(\textbf{H}_{i,d}\textbf{W}_{b}+c)$ (19)
@@
IV-D Fine-Grained Movement Prediction